Compute the IRR statistic for Project F. The appropriate cost of capital is 12 percent. (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Project F
Time: 0 1 2 3 4
Cash flow: −$10,300 $3,950 $4,780 $2,120 $2,750



Should the project be accepted or rejected?

  • rejected

  • accepted
Answer #1

Net present cash outflow = 10300

Net present cash inflow = cash inflow /1+coc x 1+coc (multiply till no. Of years it takes to generate inflow)

Year 1. 3950/(1.12) = 3537

Year 2. 4780/(1.12x1.12). = 3810

Year. 3. 2120/(1.12x1.12x1.12). = 1509

Year 4. 2750/(1.12x1.12x1.12x1.12). = 1747

Total net cash inflow. = 10603

Net present value = net cash inflow - net cash outflow

= 10603. - 10300

=. 300

As net present value is positive the project shall accepted.